OpenStudy (anonymous):

FAN N MEDAL What is the volume of a cone (in cubic inches) with a radius of 6 inches and a height of 3 inches? Use 3.14 for π. Round your answer to the nearest hundredth. 37.68 cubic inches 56.52 cubic inches 100.48 cubic inches 113.04 cubic inches

OpenStudy (anonymous):

is it a right, circular-based cone?

OpenStudy (calculusxy):

the volume of a cone is 1/3 of the volume of a cylinder a volume of cylinder would be area of base * height \[B \times height = (\pi r^2 ) \times height = 3.14(6)^2 \times 3 = 339.12 \rightarrow 339.12 \div 3 = ? \]

OpenStudy (anonymous):

113.04

OpenStudy (anonymous):

i think

OpenStudy (calculusxy):

correct!

OpenStudy (anonymous):

thank you

OpenStudy (calculusxy):

you're welcome !